### v1

# Low-energy threshold demonstration for dark matter searches in TREX-DM with an 37Ar source produced at CNA HiSPANoS

Abstract
We report on the successful implementation of an 37Ar calibration source in the TREX-DM detector, a high-pressure time projection chamber designed for low-mass dark matter searches. The 37Ar source was produced through fast neutron activation of CaO powder at the HiSPANoS facility of Centro Nacional de Aceleradores (CNA) in Spain, yielding kBq of activity. Using a novel combined GEM-Micromegas readout system, we successfully detected both characteristic emissions from 37Ar decay (2.82 keV and 270 eV) and achieved unprecedented energy threshold performance in TREX-DM, approaching the single-electron ionization energy of argon.
1 Introduction
Dark matter is a form of non-luminous, non-baryonic matter that makes up approximately 26% of the mass-energy budget of the Universe [1]. Among all the particles proposed to constitute dark matter, Weakly Interacting Massive Particles (WIMPs) stand out as one of the leading candidates. However, the traditional mass range motivated by the so-called ”WIMP miracle” [2] has been increasingly constrained by numerous direct search experiments [3] and by collider searches at facilities such as the LHC [4]. As a result, WIMP searches are gradually shifting their focus towards the low-mass frontier ( GeV/c2), which remains the least explored region of the parameter space. This shift has been possible by significant improvements in detector sensitivity, which have allowed experiments to achieve very low energy thresholds (sub keVee).
TREX-DM [5, 6, 7] is a high-pressure time projection chamber (TPC) designed to search for low-mass WIMPs using argon- or neon-based mixtures. The selection of light noble gases is motivated by their enhanced sensitivity to low-energy nuclear recoils characteristic of WIMP interactions, while operation at high pressure increases the target mass density and detection efficiency.
The detector consists of a cylindrical vessel made of radiopure copper, with a diameter of 0.5 m, length of 0.5 m, and wall thickness of 6 cm. These dimensions are determined by the requirements of holding pressures up to 10 bar(a) while serving as the innermost component of the shielding system. The vessel houses a total active volume of 20 L, corresponding to approximately 0.3 kg of argon or 0.16 kg of neon at 10 bar.
A central cathode, connected to high voltage through a custom-made feedthrough, divides the vessel into two symmetric sensitive volumes. Each volume includes a 16-cm-long drift region defined by a set of copper strips printed on a Kapton substrate and supported by four PTFE walls. This configuration ensures uniform electric field throughout the active volume.
The readout system is composed of two cm2 Microbulk Micromegas detectors [8, 9], placed at opposite ends of the active volumes. Each detector is segmented into 256 channels in the X direction and 256 in the Y direction, with a strip pitch of approximately 1 mm. This layout provides excellent spatial resolution [10] and background discrimination capabilities. Micromegas technology is particularly well-suited for rare event searches due to its intrinsically low radioactivity [11] and potential for achieving low-energy thresholds [12]. Furthermore, recent work has shown that stacking a GEM preamplification stage on top of the microbulk Micromegas, such as the ones used in TREX-DM, can provide extra preamplification factors in the range [13], contingent on gas pressure.
The experiment is located at the Laboratorio Subterráneo de Canfranc (LSC) under the Spanish Pyrenees, where the rock overburden provides natural shielding against cosmic muons and associated secondary particles.
The energy threshold and resolution directly impact the sensitivity reach for low-mass WIMP searches, where nuclear recoil energies can be as low as a few keV or below [14]. Understanding detector response and achieving precise event reconstruction at such low energies is critical for distinguishing potential WIMP signals from background events. Most commonly used external calibration sources suffer from detector self-shielding and limited penetration depth, resulting in non-uniform event distributions throughout the active volume. Furthermore, most standard calibration sources lack discrete emissions in the sub-keV energy range. For this purpose, 37Ar has emerged as an ideal calibration source. This gas radioisotope ( d) decays to 37Cl via electron capture ( keV), creating vacancies in the atomic electron shells. The subsequent atomic relaxation process produces discrete, monoenergetic emissions: 2.82 keV from K-shell capture (90% probability) and 0.27 keV from L-shell capture (9% probability) [15]. The gaseous nature of 37Ar ensures a homogeneous illumination across the sensing planes, eliminating spatial biases inherent to external point sources.
This work presents the development and implementation of a 37Ar calibration source for the TREX-DM detector, demonstrating its capability to achieve exceptionally low energy thresholds. In Section 2, we outline the chosen production method and the irradiation process carried out at Centro Nacional de Aceleradores (CNA) in Sevilla. Section 3 analyzes the low-energy calibration spectrum obtained with the TREX-DM detector, where an exceptional energy threshold of eV was achieved thanks to the gain provided by the combined GEM-Micromegas structure. Finally, Section 4 summarizes our findings and discusses their implications for low-mass WIMP searches.
2 Production Method and Irradiation at CNA HiSPANoS
2.1 Production Method and Setup
Several production routes have been investigated for 37Ar generation:
Activation of 36Ar with thermal neutrons via the reaction 36Ar(n,)37Ar. This can be achieved using atmospheric argon [18] or, as implemented by the XENON collaboration, using argon enriched in 36Ar [19]. The enrichment minimizes the production of unwanted isotopes such as 39Ar ( years) and 41Ar ( minutes), while maximizing 37Ar yield.
Neutron activation of calcium oxide (CaO) via 40Ca(n,)37Ar. In this case, the use of thermal neutrons is possible, but the cross section for fast neutrons dominates the reaction [20]. This represents the most common method for 37Ar production, and it has demonstrated success across multiple experimental contexts, including dual-phase xenon TPCs [21] and gas-based TPCs such as the one used by NEWS-G [22].
For implementation in TREX-DM, the CaO irradiation method was selected based on the ready availability of CaO powder and the possibility to access a suitable fast neutron facility, thanks to our collaboration with CNA.
The set-up for irradiation comprised a stainless-steel cross-shaped vessel (CF DN40 - CF DN100) filled with 0.5 kg of high-purity CaO powder (99.9% purity from Sigma Aldrich). Two 0.5 m particulate filters (Swagelok) were installed on the vessel sides to prevent powder contamination during gas transfer operations. An Ultra High Vacuum (UHV) valve mounted on the top part of the cross ensures appropriate vacuum levels can be achieved and maintained, and a 4 m filter is placed in the UHV line to protect the turbopump during evacuation. The complete system underwent comprehensive leak testing up to 10 bar. A custom support structure was fabricated to ensure mechanical stability during transportation between facilities and handling operations. An image of the set-up can be seen in Figure 1.
2.2 Irradiation Process at CNA HiSPANoS
The neutron irradiation was conducted at the Centro Nacional de Aceleradores (CNA) in Sevilla using their HiSPANoS neutron beam facility [23, 24]. At HiSPANoS, continuous or pulsed beams of thermal, epithermal or fast neutrons can be produced through the interaction of proton or deuteron beams of up to 6 MeV with D, Li or Be targets. For the purpose of efficient production of 37Ar, the neutron yield and energy needed to be maximized. Hence, n/sr/s fast neutrons were produced bombarding an air-cooled beryllium target with a 5 MeV deuteron beam at 1.2 A. The irradiation consisted of approximately 7 h of beam exposure, with the CaO vessel positioned at 75 mm from the beryllium target (see Figure 2).
In addition to 37Ar, natural calcium irradiation simultaneously produces 42K via the reaction 42Ca(n,p)42K, which serves as a valuable production monitor through its characteristic 1525 keV gamma emission ( h). Although the natural abundance of 42Ca (0.647%) is significantly lower than 40Ca (96.941%), the reaction still provides practical means to assess reaction success when direct 37Ar detection is not feasible. To track this gamma emission, two scintillation detectors (NaI and LaBr3) were used, previously calibrated with standard sources (137Cs at 662 keV, 60Co at 1173 and 1332 keV). The extrapolated 37Ar activity was found to be kBq.
Following neutron irradiation, the vessel radioactivity was monitored using a dosimeter until ambient background levels were achieved ( Sv/h), which required 4-5 days. The source was then transported directly to LSC, with a total time of two weeks between irradiation and injection into TREX-DM. This delay resulted in approximately 25% activity loss due to radioactive decay, a negligible amount given the initial kBq-level production.
3 Calibration in TREX-DM
The 37Ar calibration measurement was conducted with the TREX-DM detector equipped with a fully operational GEM-Micromegas readout system like the one presented in our previous work [13]. The gas mixture used in this calibration is Ar-1%iC4H10 and the detector pressure 1 bar. To process and analyze the data, we used REST-for-Physics [25], a ROOT-based software framework developed for the analysis of data from rare event search experiments.
The injection of 37Ar into the TREX-DM chamber was performed two weeks after irradiation through a controlled multi-cycle process. The irradiated container was first filled with Ar-1%iC4H10 gas mixture at 2 bar, then connected to the detector gas system maintained at 0.9 bar. After pressure equilibration, the chamber was isolated and the container refilled. This procedure was repeated four times, with each cycle extracting approximately 50% of the remaining 37Ar activity and increasing chamber pressure by 0.04 bar per cycle.
3.1 Low-Energy Spectrum
Once the source was injected into the detector volume, data acquisition was performed at high-voltage settings ( V, V) to maximize low-energy sensitivity.
Figure 3 presents the complete 37Ar energy spectrum obtained with TREX-DM. The blue curve corresponds to the full GEM-MM configuration with both amplification stages active, while the red curve shows the spectrum acquired with only the Micromegas stage operational ( V). The difference between these configurations demonstrates the significant preamplification provided by the GEM foil. The spectrum clearly reveals both characteristic 37Ar emissions: the 2.82 keV peak from K-shell electron capture and, remarkably, the 270 eV peak from L-shell capture. The observation of this ultra-low-energy peak represents a significant achievement for TREX-DM, demonstrating sensitivity in the energy region critical for light WIMP searches.
The low-energy portion of the spectrum (right panel of Figure 3) shows that events are detected down to eV, approaching the single-electron ionization energy of argon. This threshold performance validates the suitability of the GEM-MM readout technology for rare event search applications.
Analysis of the additional preamplification factor provided by the GEM shows close agreement with test bench measurements: comparing the 2.82 keV peak position in the Micromegas-only configuration (equivalent energy 50 eV) with that of the GEM-MM configuration yields a preamplification factor of , consistent with our previous characterization studies [13].
3.2 Spatial Homogeneity Assessment
The spatial distribution of 37Ar events throughout the detector active volume was evaluated through analysis of the reconstructed 2D positions. Figure 4 shows the hitmap corresponding to the integrated calibration runs, demonstrating excellent spatial uniformity across the detector area.
Quantitative assessment of the spatial homogeneity was performed by analyzing the statistical distribution of event counts within a 2020 cm2 fiducial region, excluding potential edge effects. The bin count distribution follows the expected Poisson statistics, with the 1D projection (right part of Figure 4) showing excellent agreement with a Gaussian fit of mean and standard deviation . The extracted standard deviation matches the expected relationship , confirming the proper uniformity of the detector response. The absence of visible dead regions or response variations across the detector area is crucial, as non-uniform detector response could introduce systematic biases in background discrimination and signal identification analyses.
3.3 Energy Threshold Determination
Looking at Figure 3, we observe a population of events with energies below 100 eV. These events originate from 2.82 keV 37Ar decays occurring in the transfer region, the gap between the GEM and Micromegas stages, where ionization electrons receive amplification only from the Micromegas. Since these events experience reduced total gain compared to drift region events (which benefit from both GEM and Micromegas amplification), they appear at lower equivalent energies in the spectrum. This demonstrates that our detection capability would extend to such low-energy events if there was a drift region population at these energies, given that our electronics threshold operates well below 100 eV. Analysis of representative low-energy events (illustrated in Figure 5) shows these signals represent genuine physical events well above baseline noise levels, with reliable event identification reaching equivalent energies as low as 30 eV.
The determination of the detector energy threshold was performed through systematic analysis of the shape of this low-energy transfer region peak as a function of amplification voltage. The threshold manifests as a low-energy cut-off in the peak profile, which becomes less pronounced as detector gain increases. This is modelled using a cropped-peak fit consisting of a Gaussian peak multiplied by a smoothed step function:
where , and are the amplitude, mean position and standard deviation of the Gaussian, is the energy threshold, and is a parameter that regulates the smoothness of the transition in the step function, which has been implemented through the error function. All five parameters are left free in the fit. In particular, the energy threshold is the key parameter that we seek to estimate.
Figure 6 presents the threshold analysis results for four different mesh voltage settings (280, 285, 290, and 295 V), with GEM voltage fixed. Thresholds range from approximately 60 eV at 280 V to 20 eV at 295 V. The systematic decrease in energy threshold with increasing voltage is more clearly seen in Figure 7 (left side), confirming the expected detector behavior. The voltage dependence of the energy threshold follows the expected relationship , where represents the amplification gain of the system. Since , with being some constant parameter, the threshold should exhibit exponential dependence on applied voltage. Figure 7 (right side) demonstrates excellent agreement with the expected behaviour through linear fitting in semi-logarithmic coordinates .
Even considering potential systematic uncertainties from energy scale nonlinearities at ultra-low energies, this analysis demonstrates that TREX-DM is capable of achieving energy thresholds at the level of single-electron ionization (26 eV in argon).
4 Conclusions
This work demonstrates the successful implementation of a 37Ar calibration source in the TREX-DM detector, achieving exceptional low-energy threshold performance, crucial for light dark matter searches. The neutron activation of CaO powder at the HiSPANoS facility of CNA in Sevilla provides a practical and effective method for 37Ar production, yielding sufficient activity for comprehensive detector characterization.
The combined GEM-Micromegas readout system provided preamplification factors of , enabling detection of both characteristic 37Ar emissions. Most notably, the successful observation of the 270 eV L-shell peak represents a significant achievement, as this ultra-low-energy signature is typically challenging to detect in gas-based detectors. The spatial uniformity analysis confirms homogeneous detector response across the active volume, providing a sensitivity map. The systematic threshold analysis reveals remarkable detector sensitivity, with energy thresholds extending down to approximately 20 eV at optimal voltage settings. This performance approaches the fundamental limit imposed by single-electron ionization. Such thresholds have a great impact on the sensitivity projections of TREX-DM to light WIMPs, as discussed in detail in [13], potentially providing access to unexplored regions of the parameter space. Future work will focus on using this calibration protocol for background characterization studies at low energies. Additionally, efforts are underway to develop more refined calibration techniques to probe this eV range. A promising approach that has been successfully implemented in the gaseous TPC field [26] is the use of a UV laser to calibrate in the individual electron regime. This would also allow us to accurately study the detection efficiency of our combined GEM-MM detectors to ultra-low-energy events near the energy threshold region, something we could not achieve in this work due to the imprecise determination of the initial 37Ar activity.
Acknowledgments
We would like to thank the Servicio General de Apoyo a la Investigación-SAI, Universidad de Zaragoza, for their technical support, and the Micro-Pattern Technologies (MPT) workshop at CERN, where both the Micromegas and the GEM used in this article were manufactured. The authors acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC-2017-AdG IAXO+, grant agreement No. 788781), from the Agencia Estatal de Investigación (AEI) under the grant agreements PID2021-123879OB-C21 and PID2022-137268NB-C51 funded by MCIN/AEI/10.13039/501100011033/FEDER, as well as funds from “European Union NextGenerationEU/PRTR” (Planes complementarios, Programa de Astrofísica y Física de Altas Energías).